In a psychology experiment, there are nine males and seven females availble to form a research group. How many different conbinations are possible if the group is to contain two males and two females?
Answer by stanbon(75887)

In a psychology experiment, there are nine males and seven females availble to form a research group. How many different conbinations are possible if the group is to contain two males and two females?
----
Ans: 9C2 * 7C2 = 36*21 = 756
=================================
Cheers,
Stan H.
===============